OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. sin 48° cos 15° - cos 48° sin 15°

OpenStudy (raden):

sinA cos B - sin B cos A = sin(A-B)

OpenStudy (raden):

in this case, A=48 and B=15

OpenStudy (anonymous):

@RadEn so sin33?

OpenStudy (raden):

yes, that's right

OpenStudy (raden):

but if u remember the identity : sin(90-x)° = cos(x)° so, sin33° = sin(90-57)° = cos57 ° thus, u have 2 answer's